Solve: x/5x+10 + x-3/x+2 = 7/5
I have arrived at no solution, but wondered what your answer might be.

Answer by stanbon(48569)

Solve: x/5x+10 + x-3/x+2 = 7/5
Factor to get:
x/5(x+2) + (x-3)/(x+2) = 7/5
LCD = 5(x+2)

x/LCD +(x-3)5/LCD =7(x+2)/LCD
Multiply thru by LCD to get:
x+5x-15 = 7x+14
-x=29
x=-29
Cheers,
Stan H.